# The Uniform Functional Interpretation with Informative Types

### v1

@@ 9.1 Uniform restricted quantifiers
$\begin{array}[]{lcl}|\forall x\unlhd_{\tau}B(\bm{y})\,A(x)|^{\bm{a}}_{\bm{b}}&\mathrm{\,is\,}&\forall x^{\sigma}(|B(x,\bm{y})|\,\to|A(x)|^{\bm{a}}_{\bm{b}})\\[5.69054pt] |\exists x\unlhd_{\tau}B(\bm{y})A(x)|^{\bm{a}}_{\bm{W}}&\mathrm{\,is\,}&\exists x^{\sigma}(|B(x,\bm{y})|\wedge\,\forall\bm{b}\in\bm{W}|A(x)|^{\bm{a}}_{\bm{b}}).\end{array}$
@@
A second example of restricted quantifiers which are uniform consists of (unrestricted) quantifications over witness-free types (see Notation 3.5). A notable case, worth remarking explicitly, concerns quantifications over bounded types (see Notation 8.3 and Definition 8.7) within the framework of the ${\mathcal{I}}^{\mathrm{mb}}$-interpretation. In this case, as we have discussed at the end of Subsection 8.1, we get the clauses of the uniform quantification.
@@ 9.2 Semi-uniform restricted quantifiers
$\begin{array}[]{lcl}|\forall x\unlhd_{\sigma}B(\bm{y})\,A(x,\bm{y})|^{\bm{a},\bm{g}}_{\bm{b}}&\mathrm{\,is\,}&\forall x^{\sigma}(\forall\bm{v}\in\bm{g}\bm{b}\,|B(x,\bm{y})|_{\bm{v}}\,\to|A(x,\bm{y})|^{\bm{a}}_{\bm{b}})\\[5.69054pt] |\exists x\unlhd_{\sigma}B(\bm{y})A(x,\bm{y})|^{\bm{a}}_{\bm{V},\bm{W}}&\mathrm{\,is\,}&\exists x^{\sigma}(\forall\bm{v}\in\bm{V}\,|B(x,\bm{y})|_{\bm{v}}\wedge\,\forall\bm{b}\in\bm{W}|A(x,\bm{y})|^{\bm{a}}_{\bm{b}}).\end{array}$
@@
Suppose that the $B$-restricted quantifiers are semi-uniform and that $C(x,n^{\mathbb{N}},\bm{y})$ is an arbitrary formula of the source language. In this case, the following principle
$\forall x\unlhd_{\sigma}B(\bm{y})\,\exists n^{\mathbb{N}}C(x,n,\bm{y})\to\exists k^{\mathbb{N}}\forall x\unlhd_{\sigma}B(\bm{y})\,\exists n\leq k\,C(x,n,\bm{y})$ (18)
@@
$\hskip-42.67912pt\forall x(B(x,\bm{y})^{\mathcal{I}}\to\exists n({\mathcal{I}}_{\mathbb{N}}(n)\wedge C(x,n,\bm{y})^{\mathcal{I}}))\to$
$\hskip 71.13188pt\exists k\,({\mathcal{I}}_{\mathbb{N}}(k)\wedge\forall x(B(x,\bm{y})^{\mathcal{I}}\to\exists n\leq k\,C(x,n,\bm{y})^{\mathcal{I}}))$
@@
Suppose that the $B$-restricted quantifiers are semi-uniform and assume that $C(x,n^{\mathbb{N}},\bm{y})$ is a formula of the source language with no positive $\mathcal{I}$}-witnesses. In this case, the following principle
$\forall k^{\mathbb{N}}\exists x\unlhd_{\sigma}B(\bm{y})\,\forall n\leq k\,C(x,n,\bm{y})\to\exists x\unlhd_{\sigma}B(\bm{y})\,\forall n^{\mathbb{N}}C(x,n,\bm{y})$ (19)
@@
$\hskip-14.22636pt\forall k^{\mathbb{N}}({\mathcal{I}}_{\mathbb{N}}(k)\to\exists x\,(B(x,\bm{y})^{\mathcal{I}}\wedge\forall n\leq k\,C(x,n,\bm{y})^{\mathcal{I}}))\to$
$\hskip 99.58464pt\exists x\,(B(x,\bm{y})^{\mathcal{I}}\wedge\forall n^{\mathbb{N}}({\mathcal{I}}_{\mathbb{N}}(n)\to C(x,n,\bm{y})^{\mathcal{I}}))$
@@ 9.4 Analytic applications
We consider a concrete example of $B$-restricted quantifiers in $\mathbf{HA}^{\omega}$, namely, when $B(f^{\mathbb{N}\to\mathbb{N}})$ is the formula $\forall n^{\mathbb{N}}(fn\leq_{\mathbb{N}}1)$. As a matter of convenient notation, in this case we write $\forall f\leq_{\mathbb{N}\to\mathbb{N}}\!1\,A(f)$ instead of the cumbersome $\forall f\unlhd_{\mathbb{N}\to\mathbb{N}}\!B(f)\,A(f)$. By axiom (14), this is equivalent to $\forall f^{\mathbb{N}\to\mathbb{N}}(\forall n^{\mathbb{N}}(fn\leq_{\mathbb{N}}1)\to A(f))$. We use a similar notation for the existential bounded quantifier.
@@
This is a new form of uniform boundedness. It is not only because it concerns the new ${\mathcal{I}}^{\mathrm{cb}}$-interpretation, but because it holds for an intuitionistic theory which supports a form of Markov’s principle (see Theorem 9.12). This is in contrast with the realizability treatment of Kohlenbach in [29]. The above result, however, does not seem to hold for the ${\mathcal{I}}^{\mathrm{mb}}$-interpretation: see Remark 9.5 for the obstruction.
@@
where $T$ is a parameter of type $\mathbb{N}\to\mathbb{N}$ and $\overline{f}(n)$ stands for the (natural number code of the) finite sequence $f0,f1,\ldots,f(n-1)$. (This code can be taken as a standard coding of the natural number binary sequence $\mathfrak{a}(f0),\mathfrak{a}(f1),\ldots,\mathfrak{a}(f(n-1))$.) This formula has the right form for the application of Proposition 9.21 because, in this case, $C(f,n)$ is $T(\overline{f}(n))=0$. Therefore, (20) is ${\mathcal{I}}^{\mathrm{mb}}$-interpretable. Obviously, $C(f,n)$ is extensional (with respect to to $f$). Therefore, the formula
@@ 9.5 On the bounded functional interpretation
We present the ${\mathcal{I}}^{\mathrm{mb}}$-interpretation with the notation of [20]. We write $A_{{\mathcal{I}}^{\mathrm{mb}}}(\bm{a},\bm{b})$ for $|A|^{\bm{a}}_{\bm{b}}$ and write $A^{{\mathcal{I}}^{\mathrm{mb}}}$ for $\tilde{\exists}\bm{a}\tilde{\forall}\bm{b}A_{{\mathcal{I}}^{\mathrm{mb}}}(\bm{a},\bm{b})$, i.e, $\tilde{\exists}\bm{a}\tilde{\forall}\bm{b}|A|^{\bm{a}}_{\bm{b}}$. The tilde notation in the quantifiers $\tilde{\forall}$ and $\tilde{\exists}$ mean that the quantifications are over monotone elements. We get:
@@
where $\sigma$ is an arithmetical type. The above clauses come from Proposition 3.14, the simplification of the quantifier clauses discussed at the end of Subsection 8.1 and, for the disjunction clause, the discussion at the end of Subsection 4.2. Notice that elements of a monotone finite set are necessarily monotone. The reader should compare these clauses with the corresponding clauses of the bounded functional (cf. section 3 of [20]).
@@
where $s$ and $t$ are terms of the source language, $\Gamma$ is a tuple of formulas whose interpretations have no negative ${\mathcal{I}}^{\mathrm{mb}}$-witnesses and $x$ and $y$ are variables that do not appear free in the conclusion. The reader familiar with our paper [20] will recognize these axioms and the rule. Note that, by Lemma 3.17 on universal statements and the refinement (2) of Subsection 3.2, the above (universal) axioms are ${\mathcal{I}}^{\mathrm{mb}}$-interpretable and the above rule preserves ${\mathcal{I}}^{\mathrm{mb}}$-interpretability.
@@
Since Bezem’s strong majorizability relation is transitive, we can take $ta=a$. As usual, we denote the $B$-restricted quantifications by $\forall x\trianglelefteq^{B}_{\sigma}y\;(\ldots)$ and $\exists x\trianglelefteq^{B}_{\sigma}y\;(\ldots)$. Clearly, these $B$-restricted quantifications are uniform (in the sense of Definition 9.10).
@@ 10 Conclusion
We presented a flexible approach to functional interpretations. Against the background of an ur interpretation (the uniform interpretation), several choices can be made, leading to different functional interpretations. We discussed, in particular, so-called canonical interpretations of function types over two possible base arithmetical interpretations of the ground type $\mathbb{N}$. We are referring to the canonical upon the precise interpretation ${\mathcal{I}}^{\mathrm{cp}}$ and to the canonical upon the bounding interpretation ${\mathcal{I}}^{\mathrm{cb}}$. Natural questions can be raised. What kind of choice principles are realized by these interpretations? Or, more ambitiously, can we state and prove characterization theorems for ${\mathcal{I}}^{\mathrm{cp}}$ and for ${\mathcal{I}}^{\mathrm{cb}}$ (as done by Yasugi in [46] concerning Gödel’s dialectica interpretation, or by Ferreira and Oliva concerning the bounded functional interpretation [20], or by van den Berg, Briseid and Safarik concerning their interpretation of nonstandard arithmetic in [3])? We also saw that these interpretations give rise to new models of Gödel’s theory $\mathsf{T}$, namely the type-theoretic structures $\mathcal{C^{=}}$ and $\mathcal{C^{\leq}}$. Do these structures and their hereditary versions have interesting properties? How are the hereditary versions of $\mathcal{C^{=}}$ and $\mathcal{C^{\leq}}$ related to type structures based on the notions of continuous or strongly majorizable functionals, due to Kleene [27] and Kreisel [34] (see also [45]), and Bezem [5], respectively? Are these structures models of bar-recursion (see [44] and [45])?
@@
We gave a particular emphasis to an ur (uniform) interpretation that operates in the background of computationally meaningful functional interpretations. We believe that the non set-theoretical principles embodied by the uniform interpretation are worth probing because they furnish not only guidance in some proof mining analyses but also a method for extending results of the mathematical literature to new settings. Furthermore, the existence and role of these set-theoretically false principles are possibly a source of interest for the philosophy and foundations of mathematics.

### v2

@@ 9.1 Uniform restricted quantifiers
$\begin{array}[]{lcl}|\forall x\triangleleft_{\sigma}B(\bm{y})A(x)|^{\bm{a}}_{\bm{b}}&\mathrm{\,is\,}&\forall x^{\sigma}(|B(x,\bm{y})|\,\to|A(x)|^{\bm{a}}_{\bm{b}})\\[5.69054pt] |\exists x\triangleleft_{\sigma}B(\bm{y})A(x)|^{\bm{a}}_{\bm{W}}&\mathrm{\,is\,}&\exists x^{\sigma}(|B(x,\bm{y})|\wedge\,\forall\bm{b}\in\bm{W}|A(x)|^{\bm{a}}_{\bm{b}}).\end{array}$
@@
A second example of restricted quantifiers which are uniform consists of (unrestricted) quantifications over witness-free types (see Notation 3.5). A notable case, worth remarking explicitly, concerns quantifications over uniform types (see Notation 8.3 and Definition 8.7) within the framework of the ${\mathcal{I}}^{\mathrm{mb}}$-interpretation. In this case, as we have discussed at the end of Subsection 8.1, we get the clauses of the uniform quantification.
@@ 9.2 Semi-uniform restricted quantifiers
$\begin{array}[]{lcl}|\forall x\triangleleft_{\sigma}B(\bm{y})\,A(x)|^{\bm{a},\bm{g}}_{\bm{b}}&\mathrm{\,is\,}&\forall x^{\sigma}(\forall\bm{v}\in\bm{g}\bm{b}\,|B(x,\bm{y})|_{\bm{v}}\,\to|A(x)|^{\bm{a}}_{\bm{b}})\\[5.69054pt] |\exists x\triangleleft_{\sigma}B(\bm{y})A(x)|^{\bm{a}}_{\bm{V},\bm{W}}&\mathrm{\,is\,}&\exists x^{\sigma}(\forall\bm{v}\in\bm{V}\,|B(x,\bm{y})|_{\bm{v}}\wedge\,\forall\bm{b}\in\bm{W}|A(x)|^{\bm{a}}_{\bm{b}}).\end{array}$
@@
Suppose that the $B$-restricted quantifiers are semi-uniform and that $C(x,n^{\mathbb{N}})$ is an arbitrary formula of the source language. In this case, the following principle
$\forall x\triangleleft_{\sigma}B(\bm{y})\,\exists n^{\mathbb{N}}\,C(x,n)\to\exists k^{\mathbb{N}}\forall x\triangleleft_{\sigma}B(\bm{y})\,\exists n\leq k\,C(x,n)$ (18)
@@
$\hskip-42.67912pt\forall x(B(x,\bm{y})^{\mathcal{I}}\to\exists n({\mathcal{I}}_{\mathbb{N}}(n)\wedge C(x,n)^{\mathcal{I}}))\to$
$\hskip 71.13188pt\exists k\,({\mathcal{I}}_{\mathbb{N}}(k)\wedge\forall x(B(x,\bm{y})^{\mathcal{I}}\to\exists n\leq k\,C(x,n)^{\mathcal{I}}))$
@@
Suppose that the $B$-restricted quantifiers are semi-uniform and assume that $C(x,n^{\mathbb{N}})$ is a formula of the source language with no positive $\mathcal{I}$}-witnesses. In this case, the following principle
$\forall k^{\mathbb{N}}\exists x\triangleleft_{\sigma}B(\bm{y})\,\forall n\leq k\,C(x,n)\to\exists x\triangleleft_{\sigma}B(\bm{y})\,\forall n^{\mathbb{N}}C(x,n)$ (19)
@@
$\hskip-14.22636pt\forall k^{\mathbb{N}}({\mathcal{I}}_{\mathbb{N}}(k)\to\exists x\,(B(x,\bm{y})^{\mathcal{I}}\wedge\forall n\leq k\,C(x,n)^{\mathcal{I}}))\to$
$\hskip 99.58464pt\exists x\,(B(x,\bm{y})^{\mathcal{I}}\wedge\forall n^{\mathbb{N}}({\mathcal{I}}_{\mathbb{N}}(n)\to C(x,n)^{\mathcal{I}}))$
@@ 9.4 Analytic applications
We consider a concrete example of $B$-restricted quantifiers in $\mathbf{HA}^{\omega}$, namely when $B(f^{\mathbb{N}\to\mathbb{N}})$ is the formula $\forall n^{\mathbb{N}}(fn\leq_{\mathbb{N}}1)$. As a matter of convenient notation, in this case we write $\forall f\leq_{\mathbb{N}\to\mathbb{N}}\!1\,A(f)$ instead of the cumbersome $\forall f\triangleleft_{\mathbb{N}\to\mathbb{N}}\!B\,A(f)$. By axiom (14), this is equivalent to $\forall f^{\mathbb{N}\to\mathbb{N}}(\forall n^{\mathbb{N}}(fn\leq_{\mathbb{N}}1)\to A(f))$. We use a similar notation for the existential bounded quantifier.
@@
Note that Proposition 9.12 and Theorem 9.18 imply that the ${\mathcal{I}}^{\mathrm{cb}}$-interpretation is able to interpret both a form of Markov’s principle and the above form of collection. This only works because we are using the bounding interpretation at the base type $\mathbb{N}$ together with the canonical interpretation of the function types. The result, however, does not seem to hold for the ${\mathcal{I}}^{\mathrm{mb}}$-interpretation: see Remark 9.5 for the obstruction.
@@
where $T$ is a parameter of type $\mathbb{N}\to\mathbb{N}$ and $\overline{f}(n)$ stands for the (natural number code of the) finite sequence $f0,f1,\ldots,f(n-1)$. (This code can be taken as a standard coding of the natural number binary sequence $\mathfrak{a}(f0),\mathfrak{a}(f1),\ldots,\mathfrak{a}(f(n-1))$.) This formula has the right form for the application of Proposition 9.21 because, in this case, $C(f,n)$ is $T(\overline{f}(n))=0$. Therefore, (20) is ${\mathcal{I}}^{\mathrm{mb}}$-interpretable. Obviously, $C(f,n)$ is extensional (with respect to $f$). Therefore, the formula
@@ 9.5 On the bounded functional interpretation
We present the ${\mathcal{I}}^{\mathrm{mb}}$-interpretation with the notation of [20]. We write $A_{{\mathcal{I}}^{\mathrm{mb}}}(\bm{a},\bm{b})$ for $|A|^{\bm{a}}_{\bm{b}}$ and write $A^{{\mathcal{I}}^{\mathrm{mb}}}$ for $\tilde{\exists}\bm{a}\tilde{\forall}\bm{b}A_{{\mathcal{I}}^{\mathrm{mb}}}(\bm{a},\bm{b})$, i.e, $\tilde{\exists}\bm{a}\tilde{\forall}\bm{b}|A|^{\bm{a}}_{\bm{b}}$. The tilde notation in the quantifiers $\tilde{\forall}$ and $\tilde{\exists}$ mean that the quantifications are over monotone elements (cf. Definition 6.4). We get:
@@
where $\sigma$ is an arithmetical type. The above clauses come from Proposition 3.14, the simplification of the quantifier clauses discussed at the end of Subsection 8.1 and, for the disjunction clause, the discussion at the end of Subsection 4.2. Notice that elements of a monotone finite set are necessarily monotone. The reader should compare these clauses with the corresponding clauses of the bounded functional interpretation (cf. section 3 of [20]).
@@
where $s$ and $t$ are terms of the source language, $\Gamma$ is a tuple of formulas whose interpretations have no negative ${\mathcal{I}}^{\mathrm{mb}}$-witnesses, and $x$ and $y$ are variables that do not appear free in the conclusion. The reader familiar with our paper [20] will recognize these axioms and the rule. Note that, by Lemma 3.17 on universal statements and the refinement (2) of Subsection 3.2, the above (universal) axioms are ${\mathcal{I}}^{\mathrm{mb}}$-interpretable and the above rule preserves ${\mathcal{I}}^{\mathrm{mb}}$-interpretability.
@@
Since Bezem’s strong majorizability relation is transitive, we can take $ta=a$. We denote these $B$-restricted quantifications by $\forall x\trianglelefteq^{B}_{\sigma}y\;(\ldots)$ and $\exists x\trianglelefteq^{B}_{\sigma}y\;(\ldots)$, to be close to the notation of [20]. Clearly, these $B$-restricted quantifications are uniform (in the sense of Definition 9.10).
@@ 10 Conclusion
We presented a flexible approach to functional interpretations. Against the background interpretation (the uniform interpretation), several choices can be made, leading to different functional interpretations. We discussed, in particular, so-called canonical interpretations of function types over two possible base arithmetical interpretations of the ground type $\mathbb{N}$. We are referring to the canonical upon the precise interpretation ${\mathcal{I}}^{\mathrm{cp}}$ and to the canonical upon the bounding interpretation ${\mathcal{I}}^{\mathrm{cb}}$. Natural questions can be raised. What kind of choice principles are realized by these interpretations? Or, more ambitiously, can we state and prove characterization theorems for ${\mathcal{I}}^{\mathrm{cp}}$ and for ${\mathcal{I}}^{\mathrm{cb}}$ (as done by Yasugi in [46] concerning Gödel’s dialectica interpretation, or by Ferreira and Oliva concerning the bounded functional interpretation [20], or by van den Berg, Briseid and Safarik concerning their interpretation of nonstandard arithmetic in [3])? We also saw that these interpretations give rise to new models of Gödel’s theory $\mathsf{T}$, namely the type-theoretic structures $\mathcal{C^{=}}$ and $\mathcal{C^{\leq}}$. Do these structures and their hereditary versions have interesting properties? How are the hereditary versions of $\mathcal{C^{=}}$ and $\mathcal{C^{\leq}}$ related to type structures based on the notions of continuous or strongly majorizable functionals, due to Kleene [27] and Kreisel [34] (see also [45]), and Bezem [5], respectively? Are these structures models of bar-recursion (see [44] and [45])?
@@
We gave a particular emphasis to a uniform interpretation that operates in the background of computationally meaningful functional interpretations. We believe that the non set-theoretical principles embodied by the uniform interpretation are worth probing because they furnish not only guidance in some proof mining analyses but also a method for extending results of the mathematical literature to new settings. Furthermore, the existence and role of these set-theoretically false principles are possibly a source of interest for the philosophy and foundations of mathematics.
Acknowledgments. We thank Gilda Ferreira for several interesting comments and remarks on an earlier draft of the paper.